A)What are the terms missing from the geometric sequence below?
___, -11 , 22, ___, 88 , -176
B) Write a recursive equation for x n + 1.. if possible. Thanks.

Answer by rothauserc(4718)

for a geometric series, the nth term is Xn = ar^(n-1) where a is the first term and r is the common ratio
we see that a = (11/2) and r is -2 and Xn = (11/2)(-2)^(n-1)
the missing numbers are
(11/2) for n = 1 and -44 for n = 4